how can i find the linear function which has the given slope and contains the given points example
1,-2 , m = -4

Answer by london maths tutor(243)

Linear function : y = mx + c where m = -4 and c = ?
Put the point (1,-2) and m = -4 into y = m*x + c to solve for c
-2 = (-4)(1) + c
-2 = -4 + c
-2 + 4 = c
c = 2
Therefore, the linear function becomes : y = -4x + 2